OpenStudy (anonymous):

An architect is considering bidding for the design of a new museum. The cost of drawing plans and submitting a model is $19,000. The probability of being awarded the bid is 0.1, and anticipated profits are $190,000, resulting in a possible gain of this amount minus the $19,000 cost for plans and a model. Please tell me how to do the problem not just the answer. Need DIRE help. What is the expected value in this situation? Round your answer to the nearest dollar.

OpenStudy (the_fizicx99):

Mh, Possible chances of winning bids: \(\ \sf 190,000 \times 0.10 - 19,000 \implies 19,000 -19,000 = 0\) Losses: \(\ -19,000 \times 0.90 = -17,100 \) Uh, check my arithmetic to make sure

OpenStudy (anonymous):

thank you!!